# Optimizing a-DCF for Spoofing-Robust Speaker Verification

### v1

@@ Abstract
Automatic speaker verification (ASV) systems are vulnerable to spoofing attacks such as text-to-speech. In this study, we propose a novel spoofing-robust ASV back-end classifier, optimized directly for the recently introduced, architecture-agnostic detection cost function (a-DCF). We combine a-DCF and binary cross-entropy (BCE) losses to optimize the network weights, combined by a novel, straightforward detection threshold optimization technique. Experiments on the ASVspoof2019 database demonstrate considerable improvement over the baseline optimized using BCE only (from minimum a-DCF of $0.1445$ to $0.1254$), representing $13\%$ relative improvement. These initial promising results demonstrate that it is possible to adjust an ASV system to find appropriate balance across the contradicting aims of user convenience and security against adversaries.
@@ I Introduction
Automatic speaker verification (ASV) seeks to verify a claimed identity based on speech evidence [1]. Despite being a matured biometric technology, ASV is susceptible to spoofing through replay [2], text-to-speech [3] and other adversaries. To counter these risks, various countermeasures (CMs) or presentation attack detectors (PADs) [4], have been proposed for detecting spoofing. Typically CMs are developed in isolation from ASV, and the two subsystems are combined to enhance robustness of the biometric system against adversaries. The common strategies include tandem (cascade) combination [5] and fusion of detection scores [6] or embeddings [7, 8]. Besides systems constructed from combination of speaker and spoof detection subsystems, other studies consider deep end-to-end networks optimized to predict the decision directly from the speech inputs [9, 10].
@@
Whether realized by combining two subsystems (ASV and CM), or by training a monolithic end-to-end model, a key consideration is how to evaluate and optimize the complete biometric authentication system. Any such system must trade-off between misses (false rejections) and false alarms (false acceptances)—proxies for user (in)convenience and security, respectively. The so-called tandem detection cost function (t-DCF) [5] can be used to evaluate the performance of cascaded combinations of ASV and CM under a Bayes-risk framework. Nonetheless, as t-DCF requires two sets of detection scores (and two detection thresholds) corresponding to the ASV and the CM subsystems, it is inapplicable for performance assessment of other types of architectures. To address these limitations, [11] proposed recently an architecture-agnostic detection cost function (a-DCF) applicable to a wider variety of system architectures. Unlike t-DCF, a-DCF requires a single detection score and a single threshold value only. This new metric is the primary metric in one of the two main tracks in the currently running ASVspoof 5 challenge [^1] https://www.asvspoof.org/ (referred June 28, 2024).
@@
For the first time, we propose a spoofing-robust ASV back-end that is directly optimized for the new a-DCF metric (as shown in Fig. 1). As an evaluation metric based on hard error counting, the original a-DCF metric is non-differentiable and cannot be directly optimized using gradient descent schemes. Luckily, hard error counts can be “softened” (made differentiable). While direct optimization of the t-DCF was addressed in [12] through reinforcement learning, this approach is limited to the tandem architectures. By integrating a differentiable version of the a-DCF into model training, we aim to enhance the performance and robustness of speaker verification systems against spoofing attacks, ultimately contributing to the advancement of secure and reliable biometric authentication technologies. Besides integrating the a-DCF metric into model training, the threshold value is also jointly optimized during model training in the proposed method.
Figure 1: Proposed a-DCF optimization method for spoofing-robust speaker verification. The model $g(x)$ takes embeddings as input and outputs score distributions, from which error rates $\hat{P}_{\mathrm{miss}}^{\mathrm{tar}}$, $\hat{P}_{\mathrm{fa}}^{\mathrm{non}}$, and $\hat{P}_{\mathrm{fa}}^{\mathrm{spf}}$ are calculated. These error rates are then used to compute soft a-DCF, which serve as the loss function for model optimization through backpropagation, ensuring improved performance by minimizing the a-DCF.
@@ II Detection Cost Functions
Conventional detection cost functions [13] are evaluation metrics used to assess the performance of binary classification systems, such as speaker verification and anti-spoofing. They focus on balancing the trade-off between false alarms and misses, assigning specific costs to each type of error. A binary classifier can be defined as a function $g:\mathcal{X}\to\rm I\!R$ that assigns a detection score $g(x)$ for a given input $x\in\mathcal{X}$. This score then can be converted into predicted binary class label $\hat{y}\in\{0,1\}$ by comparing it with a pre-set threshold $\tau$ by $\hat{y}=\mathrm{I}(g(x)>\tau)$. Here, $\mathrm{I}$ is the indicator function which returns $1$ if the condition is true (null hypothesis accepted) and 00 otherwise. In conventional ASV, $x$ consists of a pair of enrollment and test utterances (presented typically as a pair of speaker embeddings), and the null hypothesis is that the same speaker is present in both. In our spoofing-aware ASV models, $x$ additionally contains spoof embedding extracted from the test utterance by a CM model.
@@ II-A DCF and a-DCF
Detection Cost Function (DCF) [13] is used for performance assessment of conventional ASV systems. It is defined as:
$\mathrm{DCF(\tau_{asv})}=C_{\mathrm{miss}}^{\mathrm{tar}}\cdot\pi_{\mathrm{tar}}\cdot P_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{asv}})+C_{\mathrm{fa}}^{\mathrm{non}}\cdot\pi_{\mathrm{non}}\cdot P_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{asv}}),$ (1)
where $C_{\mathrm{miss}}^{\mathrm{tar}}$ is the cost of a miss, $C_{\mathrm{fa}}^{\mathrm{non}}$ is the cost of a false alarm, $P_{\mathrm{miss}}^{\mathrm{tar}}$ is the miss rate, $P_{\mathrm{fa}}^{\mathrm{non}}$ is false alarm rate, $\pi_{\mathrm{tar}}$ is the prior probability of the target and $\pi_{\mathrm{non}}=1-\pi_{\mathrm{tar}}$ is the prior probability of nontarget, and $\tau_{\mathrm{asv}}$ is the detection threshold. The DCF in Eq. (1) allows one to set the cost parameters according to the user’s application specification.
The architecture-agnostic detection cost function (a-DCF) [11] is a performance metric introduced for spoofing-aware speaker verification (SASV) task. It extends upon the DCF in Eq. (1) by including a third class (spoofing attack):
@@
where $P_{\mathrm{fa}}^{\mathrm{spf}}$ is the false alarm for spoof trials, $\pi_{\mathrm{spf}}$ is the prior probability of spoofing, $C_{\mathrm{fa}}^{\mathrm{spf}}$ is the respective costs for spoof false alarms. $\tau_{\mathrm{sasv}}$ is the detection threshold for spoofing-aware speaker verification (SASV) system.
The miss rate ($P_{\mathrm{miss}}^{\mathrm{tar}}$) and the false alarm rates ($P_{\mathrm{fa}}^{\mathrm{non}}$, $P_{\mathrm{fa}}^{\mathrm{spf}}$) in the a-DCF metric are computed as follows:
@@
where $N_{\mathrm{tar}}$, $N_{\mathrm{non}}$, and $N_{\mathrm{spf}}$ represent the number of target, non-target, and spoof class trials, respectively.
@@ II-B Differentiable a-DCF Metric
The indicator function used to compute false alarm and miss rates in Eq. (3) is non-differentiable. Therefore, the a-DCF metric can not be directly used as loss function to optimize an SASV system. In this study, we propose a differentiable (soft) version of a-DCF suitable for direct minimization using gradient descent. Similar approaches have been successfully applied to optimize DCF for text-dependent speaker verification [14] and t-DCF for the constrained class of tandem systems [15], but not the a-DCF. The a-DCF metric provides a more holistic evaluation by integrating the costs associated with speaker verification. This approach simplifies the benchmarking process for different ASV system architectures, as it eliminates the need for separate metrics for each component. By focusing on a single metric, the a-DCF metric promotes development of integrated systems that are optimized for spoofing-aware speaker verification.
@@
where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function. Essentially, the error rates are approximated based on their distance from the threshold. The soft a-DCF loss is now defined as:
$\mathcal{L}_{\mathrm{a\!-\!DCF}}^{\mathrm{soft}}(\tau_{\mathrm{sasv}})=$ $\;C_{\mathrm{miss}}^{\mathrm{tar}}\cdot\pi_{\mathrm{tar}}\cdot\hat{P}_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{sasv}})$ (5)
$+C_{\mathrm{fa}}^{\mathrm{non}}\cdot\pi_{\mathrm{non}}\cdot\hat{P}_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{sasv}})$
$+C_{\mathrm{fa}}^{\mathrm{spf}}\cdot\pi_{\mathrm{spf}}\cdot\hat{P}_{\mathrm{fa}}^{\mathrm{spf}}(\tau_{\mathrm{sasv}}).$
@@ II-C Optimization of soft a-DCF
2:Output: Best threshold $\hat{\tau}_{*}$, Model Parameters $\Theta_{*}$
@@
16:         $\hat{\tau}\leftarrow\tau_{i+1}$
17:    end if
18:end for
19:return $\hat{\tau}_{*}$, $\Theta_{*}$
The algorithm 1 presents the training loop of a neural network towards its optimization, with both soft a-DCF ($\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(X;\Theta_{i},\tau_{i})$ ) and BCE loss ( $\mathcal{L}_{\mathrm{BCE}}(X;\Theta_{i})$ ) functions. The BCE loss is the binary cross entropy loss defined as:
$\mathcal{L}_{\mathrm{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}[y_{i}\log(\hat{y_{i}})+(1-y_{i})\log(1-\hat{y_{i}})]$ (6)
where $N$ represents the number of training samples. Essentially, the soft a-DCF loss is a loss function of application dependent a-DCF metric and useful for fine-tuning the system towards a desired miss-false alarm balance. The BCE loss, in turn, helps to separate target from the pool of the two negative classes. For these reasons, we use the combination of these two losses. The algorithm initializes a minimum a-DCF to be infinity and a fixed threshold $\tau$. Then, for each epoch, it optimizes the model weights by calculating the combined loss $\mathcal{J}(\Theta_{i},\tau_{i})$ from soft a-DCF and BCE losses and updating the weights of the neural network. This combined loss is defined as:
$J(\Theta_{i},\tau_{i})=\frac{1}{2}\bigg{(}\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(\tau_{\mathrm{sasv}})+\mathcal{L}_{\mathrm{BCE}}\bigg{)}.$ (7)
The algorithm then conducts a grid search (Line 12 in Algorithm 1) for a value of $\tau$ that gives the optimal threshold according to the a-DCF loss on the training data. The threshold value serves as the independent variable in Eq. (5) of the a-DCF loss. Using the a-DCF loss (or its combination with BCE loss) as the objective function during model training involves minimizing its value, which depends on the threshold value. This relationship necessitates finding the optimal threshold value that minimizes the a-DCF loss at each epoch which is achieved through grid search. This algorithm tracks the best model parameters and threshold as scored on development data, thus updating the lowest a-DCF found so far and its corresponding parameters. At the end, it returns the best threshold $\hat{\tau}_{*}$ and model parameters $\Theta_{*}$.
@@ III Experimental Setup, Results and Discussion
We perform experiments on the ASVspoof2019 logical access (LA)  [3] dataset. The description of the dataset and experimental details are discussed in the following subsections.
@@ III-A Dataset
The ASVspoof2019 LA subset consists of three parts, training, development, and evaluation. Each partition has a disjoint set of speakers. The average duration of the utterances is $2$ to $3$ seconds. The spoofed utterances are generated using $19$ VC and TTS algorithms ($6$ types of attacks in the training and the development sets, $13$ in the evaluation set).
@@ III-B Experimental Setup
To demonstrate our proposed optimization scheme, we consider the DNN-based embedding fusion strategy ’Baseline2’ described in [7], that was used as a baseline in the SASV2022 challenge. Thereafter, many other studies follow similar approach to perform SASV [16, 17]. We aim to show whether our proposed technique brings any performance improvement or not. This model concatenates three embeddings ($\textbf{e}=\left[\textbf{e}_{\text{enr}}^{\text{asv}},\textbf{e}_{\text{tst}}^{\text{asv}},\textbf{e}_{\text{tst}}^{\text{cm}}\right]$): two speaker (ASV) embeddings from the enrolment ($\textbf{e}_{\text{enr}}^{\text{asv}}$) and test ($\textbf{e}_{\text{tst}}^{\text{asv}}$) utterances extracted using ECAPA-TDNN [18], and one spoofing (CM) embedding ($\textbf{e}_{\text{tst}}^{\text{cm}}$) extracted from the test utterance using AASIST [19]. The concatenated embeddings are fed to a DNN model that consists of three fully connected hidden layers with leaky ReLU activation functions. The layers contain $256$, $128$, and $64$ neurons, respectively. Unlike baseline, which contains two neurons with softmax, our model’s final layer contains a single neuron with a sigmoid activation function. This is because of our a-DCF optimization requires only one output score.
We consider four back-end models that are otherwise identical but optimized differently. These methods are:
@@
  • S2: optimized using soft a-DCF loss with, $\tau_{\mathrm{sasv}}=0.5$;
  • S3: optimized using soft a-DCF loss + BCE, with $\tau_{\mathrm{sasv}}=0.5$;
  • S4: optimized using soft a-DCF loss + BCE, including threshold optimization, as detailed in Algorithm 1.
The cost values $C_{\mathrm{miss}}^{\mathrm{tar}}$, $C_{\mathrm{fa}}^{\mathrm{non}}$, and $C_{\mathrm{fa}}^{\mathrm{spf}}$ in the soft a-DCF calculation are set as $1$, $10$, and $20$, respectively, while the prior values $\pi_{\mathrm{tar}}$, $\pi_{\mathrm{non}}$, and $\pi_{\mathrm{spf}}$ are set as $0.9$, $0.05$, and $0.05$. The higher relative cost value for spoof false alarms ($C_{\mathrm{fa}}^{\mathrm{spf}}=20$) is chosen based on the importance of spoofing robustness in security-sensitive domains, such as banking, where preventing unauthorized access is crucial. This reflects the emphasis to penalize false acceptance of spoof attempts to ensure the integrity and security of such systems. Additionally, the high target prior ($\pi_{\mathrm{tar}}=0.9$) indicates that the system is designed under the assumption that most interactions are legitimate. This is practical for applications where legitimate access attempts are frequent, and it minimizes the disruption to bonafide users by reducing the likelihood of false rejections. Except training, we considered the a-DCF metric given in Eq. 2 in all the experiments for performance evaluation.

### v2

@@ Abstract
Automatic speaker verification (ASV) systems are vulnerable to spoofing attacks. We propose a spoofing-robust ASV system optimized directly for the recently introduced architecture-agnostic detection cost function (a-DCF), which allows targeting a desired trade-off between the contradicting aims of user convenience and robustness to spoofing. We combine a-DCF and binary cross-entropy (BCE) with a novel straightforward threshold optimization technique. Our results with an embedding fusion system on ASVspoof2019 data demonstrate relative improvement of $13\%$ over a system trained using BCE only (from minimum a-DCF of $0.1445$ to $0.1254$). Using an alternative non-linear score fusion approach provides relative improvement of $43\%$ (from minimum a-DCF of $0.0508$ to $0.0289$).
@@ I Introduction
Automatic speaker verification (ASV) seeks to verify a claimed identity based on speech evidence [1]. Despite being a matured biometric technology, ASV is susceptible to spoofing through replay [2], text-to-speech [3] and other attacks. To counter these risks, various countermeasures (CMs) or presentation attack detectors (PADs) [4], have been proposed for detecting spoofing. Typically CMs are developed in isolation from ASV, and the two subsystems are combined to enhance security. The common strategies include tandem (cascade) combination [5] and fusion of detection scores [6] or embeddings [7, 8]. Besides systems constructed from combination of speaker and spoof detection subsystems, monolithic end-to-end solutions [9, 10], [11] have also been considered.
@@
Whether realized as a combination of the ASV and CM subsystems, or an end-to-end model, a key consideration is how to evaluate and optimize the complete system, taking into account the trade-off between misses (false rejections) and false alarms (false acceptances)—proxies for user (in)convenience and security, respectively. The so-called tandem detection cost function (t-DCF) [5] can be used to evaluate CM-ASV cascades under a Bayes-risk framework. Nonetheless, t-DCF requires two sets of detection scores (and two thresholds) corresponding to the two subsystems, making it inapplicable for evaluation of any other type of architecture. To address this limitation, [12] proposed recently an architecture-agnostic detection cost function (a-DCF) applicable to a wider variety of architectures. Unlike t-DCF, a-DCF requires only one set of detection scores and one detection threshold. This new metric was the primary metric in one of the two main tracks in the ASVspoof 5 challenge [13].
Figure 1: The proposed a-DCF optimization for spoofing-robust speaker verification uses embeddings or scores from enrollment ($\mathcal{O}_{\text{enr}}$) and test ($\mathcal{O}_{\text{tst}}$) inputs to give scores $g(x)$ for target, non-target, and spoof classes. Error rates $\hat{P}_{\mathrm{miss}}^{\mathrm{tar}}$, $\hat{P}_{\mathrm{fa}}^{\mathrm{non}}$, and $\hat{P}_{\mathrm{fa}}^{\mathrm{spf}}$, computed using threshold $\tau_{\text{sasv}}$, form the soft a-DCF loss for backpropagation to optimize class-specific detection accuracy.
@@
For the first time, we address spoofing-robust ASV directly optimized for the a-DCF metric (Fig. 1). As a metric based on hard error counting, a-DCF is non-differentiable and hence cannot be directly optimized using gradient descent schemes. Luckily, hard error counts can be “softened” (made differentiable). While direct optimization of the t-DCF was addressed in [14] through reinforcement learning, this approach is limited to the tandem architectures. By integrating a differentiable version of the a-DCF into model training, we aim to enhance the performance and robustness of speaker verification systems against spoofing attacks, ultimately contributing to the advancement of secure and reliable biometric authentication technologies. Besides integrating the a-DCF metric into model training, the threshold ($\tau_{\mathrm{sasv}}$) is jointly optimized with the model parameters in the proposed method.
@@ II Detection Cost Functions
Detection cost functions [15] are evaluation metrics used to assess the performance of binary classifiers, such as speaker verification and anti-spoofing systems. A binary classifier can be defined as a function $g:\mathcal{X}\to\mathbb{R}$ that assigns a detection score $g(x)$ for an input $x\in\mathcal{X}$, converted subsequently into predicted label $y\in\{0,1\}$ by comparing it with a threshold $\tau$, i.e. $y=\mathrm{I}(g(x)>\tau)$. Here, $\mathrm{I}(\cdot)$ is the indicator function which returns $1$ if the condition is true (null hypothesis accepted) and 00 otherwise. In ASV, $x=(\mathcal{O}_{\text{enr}},\mathcal{O}_{\text{tst}})$ consists of a pair of enrollment $(\mathcal{O}_{\text{enr}})$ and test $(\mathcal{O}_{\text{tst}})$ utterances. The null hypothesis is that the same speaker produced both utterances. To develop a spoofing-aware speaker verification (SASV) system, we represent $x$ either as a triplet of multivariate embeddings $(\boldsymbol{\mathrm{e}}_{\text{enr}}^{\text{asv}},\boldsymbol{\mathrm{e}}_{\text{tst}}^{\text{asv}},\boldsymbol{\mathrm{e}}_{\text{tst}}^{\text{cm}})$, or as a pair of scalars $(s_{\text{asv}},s_{\text{cm}})$. In the former case, the two speaker embeddings $(\boldsymbol{\mathrm{e}}_{\text{enr}}^{\text{asv}},\boldsymbol{\mathrm{e}}_{\text{tst}}^{\text{asv}})$ extracted using an ASV model are augmented by a spoof embedding $\boldsymbol{\mathrm{e}}_{\text{tst}}^{\text{cm}}$ extracted using a CM model. The latter case, consists of a pair of ASV and CM scores We demonstrate our back-end optimization for both types of input representations.
@@ II-A DCF and a-DCF
Detection Cost Function (DCF) [15] is used for performance assessment of conventional ASV systems. It considers consequences of both false alarms (FAs) and misses through a cost model. The DCF is formally defined as,
$\footnotesize\mathrm{DCF(\tau_{asv})}=C_{\mathrm{miss}}^{\mathrm{tar}}\cdot\pi_{\mathrm{tar}}\cdot P_{\mathrm{miss}}^{\mathrm{tar}}(\tau_{\mathrm{asv}})+C_{\mathrm{fa}}^{\mathrm{non}}\cdot\pi_{\mathrm{non}}\cdot P_{\mathrm{fa}}^{\mathrm{non}}(\tau_{\mathrm{asv}}),$ (1)
where $C_{\mathrm{miss}}^{\mathrm{tar}}$ and $C_{\mathrm{fa}}^{\mathrm{non}}$ are the costs of a miss and FA, respectively, $P_{\mathrm{miss}}^{\mathrm{tar}}$ and, $P_{\mathrm{fa}}^{\mathrm{non}}$ are the miss and FA rates, $\pi_{\mathrm{tar}}$ and $\pi_{\mathrm{non}}=1-\pi_{\mathrm{tar}}$ represent target and nontarget prior probabilities, and $\tau_{\mathrm{asv}}$ is the detection threshold. The DCF in Eq. (1) allows tuning of cost parameters to specific application needs.
Architecture-agnostic detection cost function (a-DCF) [12] is used for assessing spoofing-robust ASV systems. It extends the DCF by including a third class (spoofing attack):
@@
where $P_{\mathrm{fa}}^{\mathrm{spf}}$, $\pi_{\mathrm{spf}}$, $C_{\mathrm{fa}}^{\mathrm{spf}}$ and $\tau_{\mathrm{sasv}}$ are the FA rate of spoofs, prior of spoofing, cost of spoof FA, and the threshold, respectively. The three error rates in (2) are computed as follows:
@@
where $\mathrm{tar}$, $\mathrm{non}$ and $\mathrm{spf}$ denote the sets of target, non-target, and spoof trials, respectively, with their counts denoted by $N_{\bullet}$.
@@ II-B Differentiable a-DCF Metric
The indicator function in Eq. (3) is non-differentiable. Therefore, the a-DCF metric can not be directly used as loss function in gradient search methods. In this work, we propose a differentiable (soft) version of a-DCF suitable for this purpose. Similar approaches have been successfully used to optimize DCF for text-dependent speaker verification [16] and t-DCF for tandem architectures [14]. Compared with t-DCF, applicable only to the unnecessarily restrictive class of tandem systems, a-DCF is applicable to arbitrary SASV systems. Refer to [12] for further elaboration.
@@
where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function. Essentially, the error rates are approximated based on their distance from the threshold. The soft a-DCF loss is obtained by replacing the error rate terms in Eq. (3) with the expressions from Eq. (4), applied within the original a-DCF formula in Eq. (2).
@@ II-C Optimization of soft a-DCF
2:Output: Best threshold $\tau_{\mathrm{sasv}}$, Model Parameters $\Theta_{*}$
@@
16:         $\tau_{\mathrm{sasv}}\leftarrow\tau_{i+1}$
17:         $\Theta_{*}\leftarrow\Theta$
18:    end if
19:end for
20:return $\tau_{\mathrm{sasv}}$, $\Theta_{*}$
The algorithm 1 presents the optimization loop of a generic model through a combination of soft a-DCF ($\mathcal{L}_{\mathrm{a-DCF}}^{\mathrm{soft}}(X;\Theta_{i},\tau_{i})$) and BCE ($\mathcal{L}_{\mathrm{BCE}}(X;\Theta_{i})$) losses. The former is defined as,
$\footnotesize\mathcal{L}_{\mathrm{BCE}}=-\frac{1}{N}\sum_{j=1}^{N}[y_{j}\log(g(x_{j}))+(1-y_{j})\log(1-g(x_{j}))],$ (5)
where $N$ is the number of training samples and $g(x_{j})$ is the model output (score). The algorithm initializes a minimum a-DCF to be infinity and a fixed threshold $\tau$. Then, for each epoch, it optimizes the model weights by calculating the combined loss $\mathcal{J}(\Theta_{i},\tau_{i})$ from soft a-DCF and BCE losses and updating the weights of the neural network. This combined loss is defined as Line 7 in Algorithm 1.
The algorithm then conducts a grid search (Line 12 in Algorithm 1) for a value of $\tau$ that gives the optimal threshold according to the a-DCF loss on the training data. The threshold value serves as the independent variable in Eq. (4) of the a-DCF loss. Using the a-DCF loss (or its combination with BCE loss) as the objective function during model training involves minimizing its value, which depends on the threshold value. This relationship necessitates finding the optimal threshold value that minimizes the a-DCF loss at each epoch which is achieved through grid search. This algorithm tracks the best model parameters and threshold as scored on development data, thus updating the lowest a-DCF found so far and its corresponding parameters. At the end, it returns the best threshold $\tau_{\mathrm{sasv}}$
To elaborate on the combined loss in Line 7, assume that the empirical proportion of the training data from each class (target, nontarget, spoof) are equal ($\sim\frac{1}{3}$) in each minibatch. Expanding the combined loss $\mathcal{J}(\Theta_{i},\tau_{i})$, after some algebraic manipulation, can be shown to give
$\mathcal{J}\propto$ $\frac{1}{N_{\text{tar}}}\sum_{x\in\text{tar}}\left[A\sigma\left(\tau_{\text{sasv}}-g(x)\right)-\log g(x)\right]$ (6)
$+\frac{1}{N_{\text{non}}}\sum_{x\in\text{non}}\left[B\sigma\left(g(x)-\tau_{\text{sasv}}\right)-\log(1-g(x))\right]$
$+\frac{1}{N_{\text{spf}}}\sum_{x\in\text{spf}}\left[C\sigma\left(g(x)-\tau_{\text{sasv}}\right)-\log(1-g(x))\right],$
where $A$, $B$ and $C$ are constants implied by the a-DCF parameters111 In specific, $A:=3C_{\text{tar}}^{\text{miss}}\pi_{\text{tar}}$, $B:=3C_{\text{non}}^{\text{fa}}\pi_{\text{non}}$, and $C:=3C_{\text{spf}}^{\text{fa}}\pi_{\text{spf}}$. and where $\propto$ signifies omission of unimportant scaling constants. The combined loss is balanced average of the three class-conditional loss terms, with the per-instance losses enclosed within the brackets. The relative contribution of the two types of terms depends on the a-DCF parameters. While the sigmoid (from a-DCF) and the log (from BCE) terms are tied through the score $g(x)$, $\tau_{\text{sasv}}$ is present only in the a-DCF terms—but tied across the class-conditional losses.
As we will demonstrate below, the interplay of the two losses is important in optimization. While BCE focuses on class separation (discrimination), a-DCF takes into account the costs of actual decisions by including the decision threshold. Together, BCE and a-DCF form a complementary loss function that not only establishes a robust classification boundary but also adapts the model to prioritize specific classes based on their operational relevance and associated costs.
@@ III-A Dataset
The SASV experiments are conducted on the ASVspoof2019 logical access (LA) dataset [3], which consists of training, testing, and evaluation subsets, with disjoint speaker identity. This dataset includes both bonafide and spoofed utterances, with spoofed signals generated using various voice conversion (VC) and text-to-speech (TTS) algorithms. A detailed dataset description is available at [3].
@@ III-B Experimental Setup
To demonstrate our proposed optimization scheme, we consider the DNN-based embedding fusion strategy ’Baseline2’ described in [7], used as a baseline in the SASV2022 challenge. Thereafter, many other studies follow similar approach to perform SASV [17, 18]. The two speaker embeddings $(\textbf{e}_{\text{enr}}^{\text{asv}},\textbf{e}_{\text{tst}}^{\text{asv}})$ are extracted through ECAPA-TDNN [19] and the spoof embedding ($\textbf{e}_{\text{tst}}^{\text{cm}}$) through AASIST [20]. The three concatenated embeddings are fed to a DNN model that consists of three fully connected hidden layers of $256$, $128$, and $64$ neurons with leaky ReLU activation functions. Unlike the baseline’s two-neuron softmax output, our model’s final layer has a single neuron with a sigmoid activation, as a-DCF optimization requires only one output score.
We consider four back-end models with identical architectures, optimized differently. They are:
@@
  • S2: optimized using soft a-DCF, with $\tau_{\mathrm{sasv}}=0.5$;
  • S3: optimized using soft a-DCF + BCE, with $\tau_{\mathrm{sasv}}=0.5$;
  • S4: optimized using soft a-DCF + BCE, including threshold optimization, as detailed in Algorithm 1.
In addition to these embedding fusion methods, we proposed a-DCF optimization of a trainable nonlinear score fusion method [21] (detailed below).
The cost values $C_{\mathrm{miss}}^{\mathrm{tar}}$, $C_{\mathrm{fa}}^{\mathrm{non}}$, and $C_{\mathrm{fa}}^{\mathrm{spf}}$ in the soft a-DCF calculation are set as $1$, $10$, and $20$, respectively, while the prior values $\pi_{\mathrm{tar}}$, $\pi_{\mathrm{non}}$, and $\pi_{\mathrm{spf}}$ are set as $0.9$, $0.05$, and $0.05$. The higher relative cost value for spoof false alarms ($C_{\mathrm{fa}}^{\mathrm{spf}}=20$) is chosen based on the importance of spoofing robustness in security-sensitive domains, such as banking, where preventing unauthorized access is crucial. This reflects the emphasis to penalize false acceptance of spoof attempts to ensure the integrity and security of such systems. Additionally, the high target prior ($\pi_{\mathrm{tar}}=0.9$) aligns with an assumption that most interactions are legitimate. The a-DCF metric in (2) is used for performance evaluation. Additionally, equal error rate (EER) is reported in selected cases. We use SV-EER (speaker verification EER) to assess the system’s ability to differentiate between target and non-target samples, and SPF-EER (spoofing EER) to assess ability to differentiate between bonafide targets and spoofing attacks.